OpenStudy (anonymous):

consider the vectors x1=(2,1)^T & x2=(6,3)^T in R^2 determine the length of each vector....then let x3=x1+x2 then determine length of x3

OpenStudy (jamesj):

\[ ||x_1 || = \sqrt{2^2 + 1^2} = \sqrt{5} \] You can figure out the rest now, yes?

OpenStudy (anonymous):

well i probably should have just showed you the later part becs i did the first part got square root of 5 and 45 i dont knw how to make up the x3 part

OpenStudy (anonymous):

btw thnx for responding

OpenStudy (jamesj):

x3 = x1 + x2 = (2,1) + (6,3) = (8,4).

OpenStudy (anonymous):

so why is the transpose sign written there? just to make life hard? sigh

OpenStudy (jamesj):

just to emphasize that they are vectors, not points on the plane. Technically, I should have included the transpose.

OpenStudy (jamesj):

note also \[ \sqrt{45} = \sqrt{3^2 \times 5} = 3 \sqrt{5} \]

OpenStudy (jamesj):

as these vectors are parallel, you should find the length of x3 is equal to the length of x1 and x2.

OpenStudy (jamesj):

...equal to the SUM of the lengths of x1 and x2.

OpenStudy (anonymous):